OpenStudy (xmando98):

Given P(Ec ) = 0.45, P(F) = 0.52, and P(EExam ImageF) = 0.16. Find P( E | Fc ). a) 0.8666 b) 0.3076 c) 0.5200 d) 0.8125 e) 0.7500 f) None of the above.

OpenStudy (xmando98):

Can someone please tell me how im supposed to solve this?

OpenStudy (mecharv):

What is exam image? is it union? or what?

OpenStudy (holsteremission):

I assume `Ec` means the complement of event E, so \(\mathbb P(E^C)=1-\mathbb P(E^C)\). From the definition of conditional probability, \[\mathbb P(E|F^C)=\frac{\mathbb P(E\cap F^C)}{\mathbb P(F^C)}\]Supposing "exam image" is an intersection, you can use the fact that \[\mathbb P(A\cap B)+\mathbb P(A\cap B^C)=\mathbb P(A)\]to write \[\mathbb P(E|F^C)=\frac{\mathbb P(E)-\mathbb P(E\cap F)}{\mathbb P(F^C)}\]Then, using the definition of the complement, you end up with \[\mathbb P(E|F^C)=\frac{1-\mathbb P(E^C)-\mathbb P(E\cap F)}{1-\mathbb P(F)}\] If "exam image" is an intersection, you can use the inclusion/exclusion principle which says that \[\mathbb P(A\cup B)=\mathbb P(A)+\mathbb P(B)-\mathbb P(A\cap B)\]then use this and complements where necessary to substitute the probabilities you're given.

OpenStudy (holsteremission):

Whoops, that last part should say "If 'exam image' is a union, ..."